OpenStudy (el_arrow):

find the arc length of the curve x = t^2 + 1 y = 8t^3 + 6 on the interval -2

OpenStudy (el_arrow):

@freckles for this problem do i have to find dx and dy before i start setting up the length formula?

OpenStudy (el_arrow):

hey man could you help me out with this problem real quick

OpenStudy (nincompoop):

ye

OpenStudy (el_arrow):

cool

OpenStudy (el_arrow):

so do i have to get the derivative of x and y or no

OpenStudy (el_arrow):

or how do i go about solving this problem

OpenStudy (el_arrow):

hello there?

OpenStudy (freckles):

yes you need to find dx/dt and dy/dt

OpenStudy (freckles):

you know power rule and constant rule apply those

OpenStudy (el_arrow):

whats the constant rule

OpenStudy (freckles):

derivative of a constant is zero

OpenStudy (el_arrow):

oh yeah

OpenStudy (freckles):

\[\frac{d(c)}{dt}=0 \text{ where } c \text{ is constant }\]

OpenStudy (el_arrow):

so their derivative would be x = 2t and y = 24t^2 right?

OpenStudy (freckles):

yeah

OpenStudy (freckles):

\[\int\limits_{-2}^{0}\sqrt{(\frac{dx}{dt})^2+(\frac{dy}{dt})^2} dt\]

OpenStudy (freckles):

and actually you mean dx/dt=2t and dy/dt=24t^2 *

OpenStudy (freckles):

because x=t^2+1 and y=8t^3+6

OpenStudy (el_arrow):

yeah thats what i meant

OpenStudy (el_arrow):

thank you for clearing that up